# LVLM-COUNT: Enhancing the Counting Ability of Large Vision-Language Models

### v1

@@ Appendix A Ablation Study
We examine the effect of each of the following stages in our method: i) area detection and ii) object-aware division (note that the object-aware division necessitates the inclusion of target segmentation stage). The experiments are designed to investigate the effect of each stage individually, as well as when the stages are combined in our pipeline. Additionally, we run an experiment for a case where both stages of area-detection and target segmentation are excluded. In this case, the object-aware division can not be performed. Thus, images are divided by equidistant straight lines into subimages. We give the name of naive division to such an approach. Moreover, we run another experiment where area detection is in place but the target segmentation is excluded and naive division is applied on the detected areas We run the ablation scenarios for two LVLMs: GPT-4o, and Gemini 1.5 Pro. Furthermore, the effect of super resolution at the final stage of LVLM-Count is also investigated for the case that GPT-4o is the LVLM. For the experiments, we randomly sample $4$ images from each category in the FSC-147 test set ($29$ categories) and report the performance on the resulting subset containing $116$ samples. In Table 4 we show the results of the ablation experiments.
Table 4: Ablation study for LVLM-Count on 116 samples from the FSC-147 test dataset. Columns marked with $\Delta$ show the performance difference between an entry and the base LVLM it uses. Green indicates improvement, while red represents degradation.
@@
Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ GPT-4o 18.58 - 56.84 - GPT-4o + Naive division 33.95 $\uparrow 15.37$ 88.75 $\uparrow 31.91$ GPT-4o + Area Detection + Naive division 34.82 $\uparrow 16.24$ 73.19 $\uparrow 16.35$ GPT-4o + Area detection 19.27 $\uparrow 0.69$ 70.64 $\uparrow 13.8$ GPT-4o + Object-aware division 14.92 $\downarrow 3.66$ 31.56 $\downarrow 25.28$ GPT-4o + Area detection + Object-aware divsion (equiv. to LVLM-Count) 11.03 $\downarrow 7.55$ 21.02 $\downarrow 35.82$ GPT-4o + Area detection + Object-aware divsion + Super resolution 12.30 $\downarrow 6.28$ 28.98 $\downarrow 27.86$ Gemini 1.5 Pro 22.07 - 64.34 - Gemini 1.5 Pro + Naive division 41.31 $\uparrow 19.24$ 76.56 $\uparrow 12.22$ Gemini 1.5 Pro + Area Detection + Naive division 43.71 $\uparrow 21.64$ 75.47 $\uparrow 11.13$ Gemini 1.5 Pro + Area detection 23.04 $\uparrow 0.97$ 61.94 $\downarrow 2.4$ Gemini 1.5 Pro + Object-aware division 15.39 $\downarrow 6.68$ 38.15 $\downarrow 26.19$ Gemini 1.5 Pro + area detection + Object-aware division (equiv. to LVLM-Count) 13.64 $\downarrow 8.43$ 36.56 $\downarrow 27.78$
Appendix B Bias to Small Numbers in Datasets and Performance of LVLMs
In our experiments, LVLMs (GPT-4o) are able to make correct predictions when the number of items to be counted is small, but errors increase as the ground truth number grows. Although, we cannot be certain, one likely reason for this behavior in LVLMs is that, during training, the counting questions these models encounter are heavily biased toward small numbers. As an example, in Figure 8 we show the distribution of ’How many’ questions in some well-known VQA datasets.
(a) TallyQA Dataset [Acharya et al., 2019]
(b) VQAv2 Dataset [Goyal et al., 2017]
Figure 8: Distribution of answers in two VQA datasets.
Appendix C Definition and Example of Simple and Complex Counting Tasks
Acharya et al. [2019] were among the first to formally categorize counting questions into simple and complex types. They applied a linguistic rule: first, they removed any substrings such as “…in the photo?” or “…in the image?”. Then, they used SpaCy to perform part-of-speech tagging on the remaining substring. They classified a question as simple if it contained only one noun, no adverbs, and no adjectives; otherwise, they deemed it complex. This rule classifies questions such as “How many dogs?” as simple and “How many brown dogs?” as complex. Following this rule, they built two splits for the TallyQA dataset: a simple split and a complex split. Since we sample our open-ended counting benchmark from the TallyQA simple and complex splits, we adopt the same classification criteria for this benchmark.
In Figure 9, we provide examples of a simple and a complex VQA question from TallyQA. The VQA question in Figure 9(a) is selected from the simple split of TallyQA. It is a straightforward counting question, merely asking for all instances of the animal. Figure 9(b), on the other hand, is selected from the complex split of TallyQA. In addition to counting, it requires detecting context and distinguishing between consonants and vowels.
(a) How many cows are visible?
(b) How many consonants are there on the green sign?
Figure 9: A simple (left image) and a complex (right image) question from TallyQA.
Appendix D Incorrect Examples in the FSC-147 Dataset
We observed some incorrect instances in the FSC-147 dataset. For example, see Figure 10. In these cases, the category names which are provided are incorrect. For these instances, the extensive knowledge embedded within the LVLMs employed in our approach proves to be a disadvantage. These models detect inconsistencies and provide a count of zero as the output, whereas other methods are misled by superficial similarities and mistakenly count the objects. A more thorough study is required to detect all the incorrect examples in FSC-147.
(a) Category name in FSC-147: apples, Answer in FSC-147: $182$, Ground truth: 00.
(b) Category name in FSC-147: sunglasses, Answer in FSC-147: $81$, Ground truth: 00.
Figure 10: Two erroneous examples from FSC-147. Figure 10(a) shows a number of crabapples while the category name in the FSC-147 is apples. Figure 10(b) features a number of glasses while the category name in FSC-147 is sunglasses.
Appendix E Alleviating the Problem of Inaccurate Detections at the Area Detection and Target Segmentation Stages of LVLM-Count
One of the limitations of LVLM-Count is that if the area detection fails to detect a relevant area, the objects of interest in that area will not be counted. Another limitation arises when the target segmentation fails to segment all the instances of the target category of objects. In such a scenario, non-segmented objects of interest might be cut through by division lines, causing the predicted total number to increase due to being counted multiple times. One simple, yet quite effective, approach that we use to overcome this limitation to a significant degree is setting the detection threshold of GroundingDINO, which is used in both of the mentioned stages, to an extremely low value. This significantly reduces the chance of missing an area or object of interest, although it might lead to some false positives. However, note that false positives do not harm the performance of LVLM-Count, as their only effect is that the object-aware division lines avoid cutting through them.
To demonstrate the effectiveness of this approach, we evaluate it on the Penguin dataset [Penguin Research, 2016]. The goal in this dataset is to count penguins in images. The challenging Penguin dataset consistently exhibits heavy occlusion and complex background patterns that can easily be mistaken for penguins Arteta et al. [2016]. This dataset consists of two splits: the mixed-site split, in which images from the same camera can appear in both the training and testing sets, and the separate-site split, in which images in each set strictly belong to different cameras. Images in this dataset are annotated by multiple annotators, where each annotator might identify a different number of penguins due to the challenges in locating them within the images. Since annotators usually undercount the penguins, similar to Arteta et al. [2016], we take the maximum number of penguins among the annotations as the ground truth and calculate MAE and RMSE with respect to this value. For more details about the dataset and the calculation of metrics, we refer the reader to Arteta et al. [2016].
Since both splits are very large, we randomly select 100 samples from each split. To preserve balance, the probability of selecting an image with a certain ground-truth annotation number is inversely proportional to the frequency of that annotation number in the entire split, excluding images with zero penguins. We set the area detection and target segmentation thresholds to 0.5, and then we run them with both thresholds set to 0.1. The results are presented in Table 5 and Table 6. We observe that for both splits, when the detection threshold is high, the MAE generally increases to a value higher than the MAE of the base LVLMs. However, setting the thresholds to a low value improves the MAE for all three LVLMs.
@@ Appendix D Incorrect Examples in the FSC-147 Dataset
Table 5: Penguin dataset - Mixed sites. Columns marked with $\Delta$ show the performance difference between LVLM-Count and the base LVLM it uses. Green indicates improvement, while red represents degradation.
@@
Method MAE (Max)$\downarrow$ $\Delta$ RMSE (Max)$\downarrow$ $\Delta$ GPT-4o 26.45 - 37.53 - LVLM-Count(GPT-4o - Thredshold=0.5) 36.06 $\uparrow$ 9.61 49.23 $\uparrow$ 11.7 LVLM-Count(GPT-4o - Thredshold=0.1) 23.54 $\downarrow$ 2.91 42.24 $\uparrow$ 4.71 Gemini 1.5 Pro 37.23 - 49.15 - LVLM-Count(Gemini 1.5 Pro - Threshold=0.5) 42.42 $\uparrow$ 5.19 50.39 $\uparrow$ 1.24 LVLM-Count(Gemini 1.5 Pro - Threshold=0.1) 20.09 $\downarrow$ 17.14 29.95 $\downarrow$ 19.2 Qwen2 VL 72B AWQ 34.25 - 55.70 - LVLM-Count(Qwen2 VL 72B AWQ - Threshold=0.5) 37.13 $\uparrow$ 2.87 53.71 $\downarrow$ 1.99 LVLM-Count(Qwen2 VL 72B AWQ - Threshold=0.1) 18.22 $\downarrow$ 16.03 28.76 $\downarrow$ 26.94
Table 6: Penguin dataset - Separate sites. Columns marked with $\Delta$ show the performance difference between LVLM-Count and the base LVLM it uses. Green indicates improvement, while red represents degradation.
Method MAE (Max)$\downarrow$ $\Delta$ RMSE (Max)$\downarrow$ $\Delta$ GPT4o 36.01 - 46.64 - LVLM-Count(GPT-4o - Thredshold=0.5) 43.25 $\uparrow$ 7.24 52.17 $\uparrow$ 5.53 LVLM-Count(GPT-4o - Thredshold=0.1) 27.11 $\downarrow$ 8.9 36.72 $\downarrow$ 9.92 Gemini 1.5 Pro 46.67 - 57.26 - LVLM-Count(Gemini 1.5 Pro - Threshold=0.5) 33.59 $\uparrow$ 13.08 46.23 $\uparrow$ 11.03 LVLM-Count(Gemini 1.5 Pro - Threshold=0.1) 28.50 $\downarrow$ 18.17 41.39 $\downarrow$ 15.87 Qwen2 VL 72B AWQ 44.96 - 66.78 - LVLM-Count(Qwen2 VL 72B AWQ - Threshold=0.5) 51.17 $\uparrow$ 6.21 75.49 $\uparrow$ 8.71 LVLM-Count(Qwen2 VL 72B AWQ - Threshold=0.1) 24.83 $\downarrow$ 20.13 38.92 $\downarrow$ 27.86
In order to also provide a visual insight into this limitation, we chose an image from the dataset where distinguishing penguins from the background features is challenging for GroundingDINO. Figure 11 shows the performance of our method on this image, with the area detection and target segmentation thresholds both set to 0.5. We can observe that the failure of the target segmentation stage in segmenting all instances of penguins in Figure 11(a) has resulted in division paths cutting through several penguins inside the areas surrounded by yellow shapes in Figure 11(b). Consequently, these penguins are counted twice, causing a larger error.
On the other hand, Figure 12 shows the performance of LVLM-Count when the area detection and target segmentation thresholds are set to 0.1. Figure 12(a) illustrates that, while a low threshold has helped to segment all instances of the penguins, it has also caused some false positive areas to be segmented. Nonetheless, it can be observed in Figure 12(b) that the only effect of false positive segmentation masks is that the division paths avoid cutting through those regions as well. Note that for both images, the cluster-based approach was used to automatically find the start and end points of the division paths based on the arrangement of the segmentation masks in the scene. Comparing Figure 11 and Figure 12, we can clearly see that reducing the detection threshold is a quite successful technique to overcome limitations in the initial stages such as target segmentation.
(a) With a threshold of 0.5, target segmentation stage has failed to segment all penguins
(b) Areas surrounded by yellow shapes contain several penguins bisected by the dividing lines.
Figure 11: In this figure, the detection threshold in the target segmentation stage is set to 0.5. For this example, the cluster-based approach has been used to find the start and end points of the division paths. Ground truth (Max) = 106, LVLM-Count prediction = 131.
(a) With a low threshold such as 0.1, some false positive areas have also been segmented.
(b) The division paths avoid cutting through the false positive areas as well.
Figure 12: In this figure, the detection threshold in the target segmentation stage is set to 0.1. For this example, a cluster-based approach was used to find the start and end points of the division paths. Ground truth (Max) = 106, LVLM-Count prediction = 110
Additionally, Figure 13 shows how using a low detection threshold can help alleviate the limitations in the area detection stage for a challenging image taken in unfavorable weather conditions. Figure 13(a) shows the output of the area detection stage when the threshold is set to 0.5. The output contains two detected areas that, in combination, fail to cover all the penguins, resulting in error. However, Figure 13(b) shows the output of the area detection stage when the threshold is set to 0.1, successfully giving the area containing all the penguins.
(a) The area detection threshold set to 0.5
(b) The area detection threshold set to 0.1
Figure 13: This figure shows that setting the detection threshold in the area detection stage to a very low value (Figure 13(b)) significantly alleviates the area detection limitations at the area detection stage.
Figure 14: Illustration of the ability of LVLM-Count in counting an object of interest determined my a prompt when multiple object categories exist in a single image. Top row: Object of interest is “person”. Middle row: Object of interest is “cow”. Bottom row: Object of interest is “person and cow”
Appendix G Real-world Application of LVLM-Count
@@ Appendix H Performance Analysis of LVLM-Count for Different Ground Truth Ranges on FSC-147 Dataset
To further investigate the performance of our pipeline, we divide the ground truth values in the FSC-147 test set into intervals and plot the MAE for the base GPT-4o and Gemini 1.5 Pro models, alongside the results from LVLM-Count using each model, as shown in Figure 18. The first interval contains relatively small ground truth values, a range where LVLMs already perform well. As the ground truth values increase, the base models exhibit increasingly larger errors compared to LVLM-Count, with the margin growing rapidly. This behavior is consistent with our observations of counting errors on the blue circles in Figure 2.
@@
(b) LVLM-Count using Gemini 1.5 Pro

### v3

@@ Appendix A Ablation study
We examine the effect of each stage in our method: (i) area detection and (ii) object-aware division using masks produced during target segmentation. We design different experiments to investigate each stage’s individual contribution, as well as their combined effect in our pipeline.
To demonstrate our method’s minimal dependence on detection accuracy in the area detection and target segmentation stages, we conduct an experiment without area detection and without using GroundingDINO for segmentation masks. Instead, we configure SAM in “segment anything” mode, which generates masks for all entities in the scene. The results show that our method remains effective even without any detection model, confirming that LVLM-Count has minimal dependence on the accuracy of these initial stages.
Additionally, we conduct an experiment where both area detection and segmentation stages are excluded. In this case, object-aware division cannot be performed, so we divide images using equidistant straight lines into subimages (which we call “naive division”). We also run another experiment with area detection but without segmentation, applying naive division to the detected areas. The results demonstrate that unlike object-aware division, the naive approach is unreliable due to repetitive counting of fragmented objects.
We perform ablation studies using GPT-4o, conducting experiments on the FSC-147 test set. Table 5 presents the results of these ablation experiments.
@@
Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ GPT-4o 25.17 - 137.86 - GPT-4o + Naive division 33.04 $\uparrow$ 7.87 116.83 $\downarrow$ 21.03 GPT-4o + Area Detection + Naive division 32.69 $\uparrow$ 7.52 102.41 $\downarrow$ 35.45 GPT-4o + Area detection 23.08 $\downarrow$ 2.09 120.06 $\downarrow$ 17.8 GPT-4o + Object-aware division (using SAM without GroundingDINO) 21.01 $\downarrow$ 4.16 135.04 $\downarrow$ 2.82 GPT-4o + Object-aware division 19.17 $\downarrow$ 6.00 120.61 $\downarrow$ 17.25 GPT-4o + Area detection + Object-aware division, (equiv. to LVLM-Count) 18.09 $\downarrow$ 7.08 88.77 $\downarrow$ 49.09
@@
Table 5: An ablation study of LVLM-Count on the FSC-147 test dataset. Columns marked with $\Delta$ show the performance difference between LVLM-Count and the base LVLM. Green indicates improvement, while red represents degradation.
Appendix B Visual Examples of LVLM-Count’s Performance on the FSC-147 Dataset, PASCAL VOC, and Emoji-Count Benchmarks
This section presents several visual examples showcasing the performance of LVLM-Count on the FSC-147 dataset , PASCAL VOC, and the Emoji-Count benchmarks. The LVLM used in the pipeline to generate these visual examples is GPT-4o.
Figure 8 illustrates three examples of LVLM-Count’s performance on FSC-147. Additionally, Figure 9 includes an example from the PASCAL VOC benchmark. Finally, we present three visual examples from the Emoji-Count benchmark in Figure 10. In these visual examples, LVLM-Count achieves more accurate results compared to the base GPT-4o.
Figure 8: Illustration of three examples of the performance of LVLM-Count on FSC-147. Top row: The object of interest is “strawberry”. Middle row: The object of interest is “hot air balloon”. Bottom row: The object of interest is “sheep”.
Figure 9: An example of the performance of LVLM-Count on the PASCAL VOC benchmark. The object of interest is “bottle”.
Figure 10: Three examples of the performance of LVLM-Count on the Emoji-Count benchmark. Top row: The object of interest is “right arrow curving left”. Middle row: The object of interest is “woman with red hair”. Bottom row: The object of interest is “hibiscus”.
Appendix C Robustness to Crowded Scenes and Heavy Occlusion
One of the key strengths of our approach is its robustness in crowded scenes and images with heavy occlusions. To achieve this robustness, the masks produced by SAM are not directly used for object-aware division. Instead, several important post-processing operations are applied beforehand. The most critical of these is non-maximum suppression, which removes lower-confidence masks that overlap with others. Additionally, an erosion function is applied to trim the outermost layer of each mask, ensuring at least two pixels of empty space between adjacent masks to allow for reliable placement of division lines. A polishing step further refines the masks by smoothing their surfaces and eliminating artifacts, preventing division lines from mistakenly passing through them.
Figure 11 demonstrates the strong performance of our method on a randomly selected image from the web containing various types of flowers. This scene features extreme occlusion; nonetheless, our method effectively handles the division task. It is important to note that the strong performance of our method in this challenging scenario is not coincidental. Figure 12 illustrates the impact of our post-processing operations. The upper row displays the unprocessed SAM masks, while the bottom row shows the results after post-processing, which generates reliable paths for the division lines.
Figure 11: An example of the performance of LVLM-Count on a random crowded image from the web, involving heavy occlusion. The object of interest is “flower”.
Figure 12: The effect of post-processing on the masks produced by SAM in making LVLM-Count robust to crowded scenes with heavy occlusions. The object of interest is “flower”.
To further demonstrate the effectiveness of our method in crowded scenes, we present Table 6, which compares the performance of different LVLMs on FSC-147 dataset samples where the ground truth count is 100 or higher. The table contrasts their baseline performance with their performance when using our LVLM-Count method. The results show significant improvements when LVLM-Count is employed. Furthermore, Figure 13 presents three visual examples of LVLM-Count’s performance on crowded images from FSC-147.
@@ Appendix C Robustness to Crowded Scenes and Heavy Occlusion
Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ GPT-4o 89.22 - 291.34 - LVLM-Count (GPT-4o as LVLM) 57.87 $\downarrow 31.35$ 185.29 $\downarrow 106.05$ Gemma 3 27B 125.69 - 330.38 - LVLM-Count (Gemma 3 27B as LVLM) 71.68 $\downarrow 54.01$ 251.48 $\downarrow 78.9$ Qwen2 VL 72B AWQ 102.64 - 302.23 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 79.66 $\downarrow 22.98$ 289.47 $\downarrow 12.76$
Table 6: Evaluation on samples with ground truth counts equal to or higher than 100 from the test set of the FSC-147 dataset. Showing the effectiveness of our method for crowded scenes.
Figure 13: Visual examples of the performance of LVLM-Count on three crowded scenes from the FSC-147 dataset.
Appendix D Additional Information and Visual Examples from the Penguin Benchmark
To demonstrate the effectiveness of our strategy of setting a low detection threshold to overcome challenges in area detection and target segmentation stages, we evaluate it on a benchmark taken from the Penguin dataset [Penguin Research, 2016]. The goal in this dataset is to count penguins in images. This is challenging since images in the dataset consistently exhibit heavy occlusion and complex background patterns that can easily be mistaken for penguins Arteta et al. [2016]. This dataset consists of two splits: i) the mixed-site split, in which images from the same camera can appear in both the training and testing sets, and ii) the separate-site split, in which images in each set strictly belong to different cameras. Images in this dataset are annotated by multiple annotators, where each annotator might identify a different number of penguins due to the challenges in locating them within the images. Since annotators usually undercount the penguins, similar to Arteta et al. [2016], we take the maximum number of penguins among the annotations as the ground truth and calculate MAE and RMSE with respect to this value. For additional details regarding the dataset and metric calculations, we direct readers to Arteta et al. [2016].
We construct our benchmark using the separate-site split. Given the dataset’s substantial size, which comprises tens of thousands of images, we randomly select 100 samples from the chosen split. To ensure balance, the sampling probability for an image with a specific ground-truth annotation count is inversely proportional to the frequency of that count in the entire split. We evaluate two variants of our pipeline. The main variant employs GroundingDINO with a low threshold for area detection and target segmentation, while the alternative variant does not use GroundingDINO at all. In this variant, masks are obtained by configuring SAM to segment any entity in the image. The results presented in Table 4 show that both variants significantly improve the counting performance of the LVLMs, demonstrating the robustness of the initial stages in our pipeline.
Figure 14 demonstrates the segmentation masks and division paths for the two variants on a sample from the Penguin benchmark. In Figure 14 (a), GroundingDINO with a low detection threshold is employed. It correctly identifies all penguin instances, though there are some false positives. These detections are subsequently segmented by SAM. Figure 14 (c) depicts the division paths found using these masks. In Figure 14 (b), GroundingDINO is not used, and SAM is configured to segment all entities in the scene. As a result, in addition to the penguins, a large number of other objects are segmented. Nevertheless, Figure 14 (d) shows that the division paths derived from these masks successfully partition the image without cutting any penguins. Note that some excessively large masks generated by SAM have been removed due to post-processing in our pipeline.
Figure 14: Illustration of robustness in detection during the initial pipeline stages: an example from the challenging Penguin benchmark. (a) GroundingDINO with a low threshold successfully detects all penguins, though a few false positives remain. These detections are then segmented by SAM. (b) Masks produced without a detection model: Here, SAM is set to segment all entities in the image. Excessively large masks are filtered out during post-processing in our pipeline. (c) Division paths are derived using the masks generated by GroundingDINO and SAM. (d) Division paths based on SAM-only segmentation effectively partition the image without cutting through any penguins. Note that the slight difference in the area of the images on the left is due to the existence of area detection stage.
Appendix E Real-world Application of LVLM-Count
@@ Appendix F LVLM-Count’s Power in Handling Multiple Object Categories in the Same Image
Figure 18: Illustration of the ability of LVLM-Count in counting an object of interest determined by a prompt when multiple object categories exist in a single image. Top row: Object of interest is “cow”. Middle row: Object of interest is “person”. Bottom row: Object of interest is “person and cow”
Appendix G False Positive Masks at the Target Segmentation Stage
@@ Appendix H Performance Analysis of LVLM-Count for different ground truth ranges on FSC-147 dataset
To further investigate the performance of our pipeline, we divide the ground truth values in the FSC-147 test set into intervals and plot the MAE for the base GPT-4o, Qwen2, and Gemma 3 models alongside the results from LVLM-Count using each model, as shown in Figure 20. The first interval contains relatively small ground truth values, a range in which LVLMs already perform well. As the ground truth values increase, the base models exhibit increasingly larger errors compared to LVLM-Count, with the margin growing rapidly. This behavior is consistent with our observations of counting errors on the blue circles in Figure 2.
@@
(b) LVLM-Count using Gemma 3
(c) LVLM-Count using Qwen2